OpenStudy (anonymous):

Two opposite vertices of a square are given by the complex numbers z1=-3+7i and z2=9-9i. Find the other two vertices of the square and give also the equation of its circumscribed circle. Any tips are appreciated.

OpenStudy (anonymous):

I was thinking to find the mid point between the two vertices and then find the other two vertices using that mid point and different directions. I already found the mid point (3-i).

OpenStudy (anonymous):

@c0rtez  Do you know rotation ?? i.e \[\large z_4 - z_1 = \frac{z_2-z_1}{|z_2-z_1|} * |z_4-z_1|\] Since |z_4-z_1| = |z_2-z_1| (side of square are equal ) Now you will get z_4. Similarly find z_3

OpenStudy (anonymous):

oops . Sorry It should be \[\large z_4 - z_1 = \frac{z_2-z_1}{|z_2-z_1|} * |z_4-z_1| * e^{i (\-pi /2)}\]

OpenStudy (anonymous):

Typo: e^(i * pi/2)

OpenStudy (anonymous):

Ok thanks I will try to solve it now.

OpenStudy (anonymous):

Sorry typo again : e^(i * -pi/2)

OpenStudy (anonymous):

So if I got it right, I can multiply z1 by e^(i * pi/2) to get z2, and use -pi/2 for z4?

OpenStudy (anonymous):

@c0rtez, It should be \[\large z_4 - z_1 = \frac{z_2-z_1}{|z_2-z_1|} * |z_4-z_1| * e^{-i \pi /2}\] Since |z_4-z_1| = |z_2-z_1| (side of square are equal ) Now you will get z_4 from above Use the same technique and find z_3